OpenStudy (anonymous):

3. A 10-foot ladder leans against a wall with its foot braced 3 feet from wall’s base. How far up the wall does the ladder reach?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 x^2 + 3^2 = 10^2 solve for x

OpenStudy (anonymous):

√91 ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

\[\Large x = \sqrt{91}\]

jimthompson5910 (jim_thompson5910):

Which is approximately \[\Large \sqrt{91} \approx 9.53939\]

jimthompson5910 (jim_thompson5910):

So the ladder reaches roughly 9.53939 ft off the ground

OpenStudy (anonymous):

okay thank you so much!

jimthompson5910 (jim_thompson5910):

yw